a son is one-half as old as his farther. the sum of their ages is 72 years. what is the age of each?
Answer by stanbon(75887)

a son is one-half as old as his father. the sum of their ages is 72 years. what is the age of each?
----
Equations:
s = (1/2)f
s+f = 72
---
Substitute for "s" and solve for "f":
(1/2)f+f = 72
(3/2)f = 72
f = (2/3)72
f = 48
-------
Solve for "s"
s = (1/2)f
s = 24
===============
Cheers,
Stan H.